### v1

# Copenhagen Survey on Black Holes and Fundamental Physics

Abstract
The purpose of this survey is to take a snapshot of the attitudes of physicists, which may be useful to sociologists and historians of science. A total of 85 completed surveys were returned out of 151 registered participants of the “Black holes Inside and out” conference, held in Copenhagen in 2024. The survey asked questions about the nature of black holes and some of the most contentious issues in fundamental physics. A number of surprising results were found. For example, some of the leading frameworks, such the cosmological constant, cosmic inflation, or string theory, while most popular, gain less than majority of votes from the participants. The only statement that gains majority approval (by 68% of participants) was that the Big Bang means that “the universe evolved from a hot dense state”, not “an absolute beginning time”. This provides reasons for caution in describing ideas as consensus in the scientific community when a more nuanced view may be justified.
I Introduction
In August 2024, Black Holes Inside and Out conference was held with a multi-disciplinary approach to black hole physics. The conference was organized by the Niels Bohr Institute and was held at the iconic Black Diamond Building in Copenhagen, Denmark. It included leading theorists and experimentalists in astrophysics, general relativity, numerical relativity, gravitational wave astronomy, and cosmology. We considered this a good opportunity to survey experts on current controversies within physics. The purpose is to take a snapshot of the physicists’ attitudes, which may be useful to sociologists and historians of science. We also compare the results of this survey to popular reports of what physicists believe and highlight some significant discrepancies. The paper is organized as follows: We start by a brief description of our methodology in Section II. Section III focuses on topics within black hole physics, while Section IV focuses on broader issues within fundamental physics. The last section summarizes the highlights and concludes the paper.
II Methods
A list of questions was proposed by the authors and then approved by conference organisers. The survey was announced during several sessions and distributed in the lobby. A total of 85 completed surveys were returned out of 151 participants listed as attending on the conference web site. However, for some questions, a few respondents ticked more than one answer. Only questions that had single answers from a participant were included in our main results. In the appendix, we show the addition of these multiple votes makes little difference to the overall picture. Results were compiled by a third party who was not involved in the survey design.
III Black Holes
In this section we report on the results of controversial topics within the area of black hole physics.
III.1 Information Paradox
Since the 1970’s it has been argued that black holes in general relativity destroy information, potentially violating the unitarity principle of quantum mechanics [8]. Several solutions to this paradox have been proposed [12] but no consensus exists. The results gathered from this survey show that most physicists (53) do believe that information is preserved either in a remnant (26), or in Hawking radiation (27). However, there is a significant minority (22) who believes that information is permanently lost in black holes, which is in tension with claims that the issue is now settled among experts, based on the holographic principle [21]. In fact, both remnants and destruction of information are in conflict with the holographic principle, as remnants have more entropy than their area permits, and information should always be preserved on the holographic boundary. The proponents of these latter two ideas add up to nearly half of the respondents (see Figure 1 and Table 1).
| Answer Option | Number of Votes | Percentage |
| It’s irretrievably lost | 17 | 22 |
| It’s preserved in Hawking Radiation | 21 | 27 |
| It’s preserved in a remnant | 20 | 26 |
| Other | 11 | 14 |
| No opinion | 8 | 10 |
| Total | 77 |
III.2 Supermassive Black Hole Formation
Supermassive black holes have been measured with masses larger than 40 billion solar masses [14] but how they form is still mysterious. Our results show that out of those that expressed an opinion, direct collapse (25) was the most popular option with primordial black holes (22) being a close second (see Figure 2 and Table 2).
| Answer Option | Number of Votes | Percentage |
| Primordial Black Holes | 17 | 22 |
| Direct collapse seeds | 19 | 25 |
| Super Eddington accretion into stellar seeds | 8 | 10 |
| Sub-Eddington accretion into stellar seeds | 5 | 6 |
| Other | 5 | 6 |
| No opinion | 23 | 30 |
| Total | 77 |
III.3 Fate of Matter in Black Holes
In classical relativity, it was proved that matter that falls into a black hole will be crushed into a singularity [17]. But do physicists believe this really happens; or are they happy to entertain more exotic alternatives? The classical view remains favoured by a substantial portion of respondents (29, see Figure 3 and Table 3). While no single alternative to the singularity garners a higher level of approval, the combined backing for the three proposed exotic solutions (38) exceeds that of the classical option.
| Answer Option | Number of Votes | Percentage |
| It is crushed into a singularity | 22 | 29 |
| It becomes a fuzzball | 12 | 16 |
| It bounces out into our universe | 9 | 12 |
| It bounces out into another universe | 8 | 10 |
| Other | 20 | 26 |
| No Opinion | 6 | 8 |
| Total | 77 |
III.4 Testing Quantum Gravity Models with Black Holes
The energy scale of quantum gravity is far beyond the reach of any particle detector on Earth. However, it has been shown that a black hole can accelerate particles to arbitrarily high levels [1]. This provides hope that black holes could be used as probes of quantum gravity. However, there is more than one possible way to use black holes for this purpose. In our survey, we see that among those who selected a particular method, Effective Field Theory corrections to inspiral/ringdown was by far the most popular view (32, see Figure 4 and Table 4). However, the same number of respondents suggested that there could be other methods to test quantum gravity with black holes, not listed in our survey. The nature of such methods remains an enigma to the authors of this article.
| Answer Option | Number of Votes | Percentage |
| No opinion | 8 | 11 |
| Other | 23 | 32 |
| Deviations from Kerr Spacetime Multipoles | 8 | 11 |
| Gravitational Wave Echoes | 5 | 7 |
| Effective Field Theory corrections to inspiral/ringdown | 23 | 32 |
| Black hole shadows are not dark | 6 | 8 |
| Total | 73 |
IV Fundamental Physics
In this section we report on results from questions concerning a broader range of controversies within physics not specifically linked to black holes.
IV.1 Interpretation of Quantum Mechanics
Quantum mechanics has been described as both a great success in physics but also “a scandal" [24] in that despite a remarkable predictive accuracy, its physical interpretation remains elusive. Our survey shows the two most well-known “interpretations" of quantum mechanics, Copenhagen and Many Worlds are indeed the most popular choices amongst those who chose a specific option, with Copenhagen remaining the preferred option (see Figure 5 and Table 5). Copenhagen’s dominance is consistent with previous surveys [13, 22]).
| Answer Option | Number of Votes | Percentage |
| Copenhagen | 23 | 28 |
| Many Worlds | 9 | 11 |
| Qbism | 0 | 0 |
| Pilot Wave/de Broglie-Bohm | 3 | 4 |
| Collapse Theories | 6 | 7 |
| Consistent Histories | 3 | 4 |
| Other | 10 | 12 |
| No Opinion | 27 | 33 |
| Total | 81 |
IV.2 Quantum Gravity
A theory of quantum gravity has been described as the “holy grail" of physics [19] but no candidate has been widely accepted by the scientific community. String theory, perhaps the most prominent proposal, has been described by commentators as anything from “the only idea about quantum gravity with any substance” [10] to a theory that is “dead” [16]. Participants in our survey who expressed a preference clearly favour string theory over any other candidate, however it is very far from consensus and some researchers believe gravity is not quantum as has been explored recently [15]. Another surprising result is that asymptotically safe quantum gravity was the second most popular theory even though a google scholar search for the third most popular candidate (loop quantum gravity) returned more than ten times the results than for asymptotic safety. Interestingly, this is the only question where those picking none of the presented physical options (combining no opinion and other) represented the majority view (51, see Figure 6 and Table 6).
| Answer Option | Number of Votes | Percentage |
| String/M Theory | 16 | 21 |
| Loop Quantum Gravity | 5 | 6 |
| Causal Sets | 0 | 0 |
| Hořava–Lifshitz gravity | 4 | 5 |
| Causal Dynamical Triangulations | 0 | 0 |
| Asymptotic Safety | 6 | 8 |
| Gravity is not quantum | 7 | 9 |
| Other | 21 | 27 |
| No opinion | 19 | 24 |
| Total | 78 | 100 |
IV.3 Dark Matter
The standard model of cosmology, LCDM, assumes most of the matter in the universe is “dark” in that it does not interact with electromagnetic radiation [23]. However, establishing which of the many dark matter candidates is part of nature (and which are not) has not been accomplished by science. According to Wikipedia “The most prevalent explanation is that dark matter is some as-yet-undiscovered subatomic particle” [25]. However, in our survey, the most popular specific candidate is “primordial black holes”, which had more votes than WIMPs and axions combined (although in the appendix we see when multiple votes are counted WIMPS do have a significant boost). Nevertheless, the fact that this survey was performed at a black hole conference may be a contributing factor to this result. Modified gravity theories remain fringe, although presumably the hybrid option could include some form of modification to gravity as has been suggested, e.g., in the superfluid model [2]. The hybrid explanation is the most popular (31, see Figure 7 and Table 7) of all options for those that expressed an opinion. A hybrid might also refer to a combination of heavy and light particles, primordial black holes and some modification to gravity. Dark matter may be ubiquitous, and this seems a popular view amongst our respondents.
| Answer Option | Number of Votes | Percentage |
| Light particle dark matter (e.g., axions | 8 | |
| Heavy particle dark matter (e.g., WIMPs) | 3 | |
| Some form of modified gravity | 1 | |
| Primordial black holes | 12 | |
| Some hybrid of the above | 22 | |
| Effect of quantum gravity | 3 | 4 |
| Other | 10 | |
| No opinion | 11 | |
| Total | 70 | 100 |
IV.4 Early Universe Cosmology
The standard picture of the Big Bang faces several difficult puzzles such as the horizon, monopole, flatness and the origin of structure problem [7]. Inflation can solve these problems and has been called the third pillar of the standard LCDM model [23]. However, our survey shows that whilst it is by far the most popular option, like string theory, it does not enjoy majority support, although it is far closer to that status than string theory is for quantum gravity (see Figure 8 and Table 8).
| Answer Option | Number of Votes | Percentage |
| Cosmic Inflation | 34 | 44 |
| Bouncing or Cyclic Universe | 6 | 8 |
| Some quantum gravity model | 12 | 16 |
| Breakdown of Lorentz Invariance | 1 | 1 |
| Not inflation but don’t have an alternative | 10 | 13 |
| Other | 5 | 6 |
| No Opinion | 9 | 12 |
| Total | 77 | 100 |
IV.5 Meaning of the Big Bang
It has been claimed that the Big Bang represents the beginning of time and this is a view now “taken for granted" [9]. However, a challenge comes from Guth who claimed that the Big Bang is “a theory of the aftermath of the bang" that says, “nothing about what banged, why it banged, or what happened before it banged" [4]. The very meaning of the term Big Bang is thus not agreed upon. Our survey results strongly favour Guth’s view over Hawking’s. The Big Bang makes no claim about whether time had a beginning or not and should be understood only as the theory that says the universe evolved from a hot dense state (68). In that sense it is perceived as a theory of what happened after the Bang only. Notably, there is more consensus on this issue than any other we surveyed. Although a counter argument to this conclusion may be that for this question there were fewer options. The classical singularity is taken far less seriously (11) than in the case of black holes (29). The number of no opinions on this topic was also the fewest (5) of any topic surveyed here (see Figure 9 and Table 9). An important implication of this result is that it may not be correct to say that scientists believe that the universe has a beginning, or even that it is 13.8 billion years old, rather it is at least 13.8 billion years old.
| Answer Option | Number of Votes | Percentage |
| An absolute beginning of time with a singularity at its start | 9 | 11 |
| An absolute beginning of time but without a singularity | 5 | 6 |
| A theory that says the universe evolved from a hot dense state that says nothing about whether there was an absolute beginning of time or not | 54 | 68 |
| Other | 8 | 10 |
| No opinion | 4 | 5 |
| Total | 80 |
IV.6 Anthropic Coincidences and Fine tuning
The constants of nature are free parameters in current physical theories. But what explains their values? Some have claimed that they are “fine-tuned for life” which requires exotic explanations either in the form of anthropic selection in a multiverse or an intelligent designer [11]. A similar question was asked of philosophers in the PhilPapers survey [3] . In the PhilPapers survey the most popular choice was the constants are considered brute facts and thus require no exotic explanation. For philosophers, that number was 32, for our survey of physicists this was also the most popular choice and the number was remarkably similar at 33. However in our case, anthropic selection in a multiverse was a clear second ( 24) whereas for philosophers this option was fourth (see Figure 10 and Table 10). One possibility that was not considered in the PhilPapers survey was a Darwinian solution to the problem (see [20] for an example). Here that option was the third most popular (10).
| Answer Option | Number of Votes | Percentage |
| The values of the constants are set by necessity | 7 | |
| They are brute facts or need no further explanation | 26 | |
| Explained by anthropic selection in a multiverse | 19 | |
| Explained by a Darwinian process occurring in the cosmos | 8 | |
| Explained by an intelligent designer | 2 | |
| Other | 7 | |
| No opinion | 10 | |
| Total | 79 |
IV.7 Dark Energy
In 1998 it was discovered that the universe is accelerating in its expansion [18]. In the standard model of cosmology this is assumed to be driven by a cosmological constant [23]. In our survey, a cosmological constant is the most popular option (38), with significantly higher support than the next most popular option, a modification to gravity at 17 (see Figure 11 and Table 11). However, it did not gain majority support, casting doubt that this is consensus amongst physicists. An important implication for science communicators is that it may not be correct to say that scientists agree the ultimate fate of the universe is a “heat death". This would be the case if there was a majority view for a cosmological constant. However, our survey shows this option falling below this threshold of support.
| Answer Option | Number of Votes | Percentage |
| A cosmological constant | 29 | 38 |
| A time varying field | 12 | 16 |
| An effect of quantum gravity | 9 | 12 |
| Some other modification to gravity | 13 | 17 |
| Other | 6 | 8 |
| No option | 8 | 10 |
| Total | 77 | 100 |
IV.8 Hubble tension
Supernova and CMB measurement can both provide a value for the Hubble parameter and should align, but they don’t [6]. This has been described by the press as a “crisis in cosmology" [5]. Our survey shows that the most popular answer by far (38, see Figure 12 and Table 12) is that the Hubble tension does not represent a sign of new physics, but rather systematic errors in supernova data.
| Answer Option | Number of Votes | Percentage |
| Systematic Error in Supernova Data | 25 | 35 |
| Systematic Error in CMB/Baryonic Acoustic Oscillation Data | 3 | 4 |
| Early Dark Energy | 4 | |
| Other New Particles or Fields | 0 | |
| Modified or Quantum Fields | 12 | |
| Other | 10 | |
| No opinion | 18 | 25 |
| Total | 72 |
V Conclusion
Polling scientists does not decide scientific controversies. However, it can be useful for sociologists and historians of science to understand how scientific views may change over time. It is also useful for communicators of science to know what leading researchers in a field think. Surveys may also be a tool to combat publication bias. For example, it may be the case that within the literature there is more discussion of new physics as a solution to the Hubble tension, than those suggesting systematic errors. But this could mask skepticism within the field about the need for exotic proposals. Counting citations might then be an incomplete probe to understand scientific opinion or gauge the extent of scientific consensus on a topic.
Any survey will have the limitation of its sample and ours is no exception. A conference on black holes might give different results to one on particle physics, foundation of physics, cosmology or solid-state physics. However, The Black Holes Inside and Out conference was multidisciplinary (although still within the realm of physics) providing some reason to think it may capture the attitudes of a wider scientific community than more specialised conferences.
Several results reveal tensions between how science is often communicated and what the scientists in our survey believe. Perhaps the biggest discrepancy is the view that The Big Bang represents the beginning of time; most physicists in our survey see the theory making much more modest claims than that. There seems to be more uncertainty for the fate of the universe as most scientists we surveyed are not convinced that the accelerated expansion is driven by a cosmological constant. Another notable discrepancy is for the claim that the “fine tuning” of the constants of nature forces us to choose between a multiverse or an intelligent designer. In our survey, brute facts that need no explanation were more popular than both; Moreover, necessity and a Darwinian explanation were more popular than an intelligent designer. In other areas, several leading candidates for solving outstanding issues in physics are the most popular option amongst scientists that expressed a preference. But in each case, they fail to garner majority support. Examples include string theory as a theory of quantum gravity, inflation as a theory of the early universe, a cosmological constant as the best explanation for the accelerated expansion, black hole information preserved as Hawking radiation, matter being crushed into a singularity inside a black hole and the Copenhagen interpretation of quantum mechanics. Particle dark matter also seems less preferred than is often communicated, although in this case a preference for primordial black holes may be biased by the selection effect of being at a black hole physics conference. We see no reason though to think this selection effect biased any of the other topics we surveyed. This survey provides reasons for caution in describing ideas as consensus in the scientific community when a more nuanced view may be justified.
Acknowledgements.
We would like to thank the organizers of the “Black Holes Inside Out” conference, especially Vitor Cardoso and Julie de Molade, for making this survey possible, and helping with distribution and collection of the questionnaires. NA and AC are funded by the University of Waterloo, the National Science and Engineering Research Council of Canada (NSERC) and the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Economic Development & Innovation.VI Appendix
Here we show results where completed surveys with multiple votes per question were included, few significant differences are seen from the main results. Two exceptions are for dark matter, where WIMPS gained a boost from 3 to 9 votes (4 to 10) and testing quantum gravity with gravitational wave echoes were boosted from 5 to 13 votes (7 to 14).
VI.0.1 Black Hole Information Loss
VI.0.2 Super Massive Black Hole Formation
| Answer Option | Number of Votes | Percentage |
| Primordial Black Holes | 20 | 22 |
| Direct collapse seeds | 23 | 25 |
| Super Eddington accretion into stellar seeds | 13 | 14 |
| Sub-Eddington accretion into stellar seeds | 7 | 8 |
| Other | 5 | 5 |
| No opinion | 23 | 25 |
| Total | 91 | 100 |
VI.0.3 Fate of Matter Inside a Black Hole
| Answer Option | Number of Votes | Percentage |
| It is crushed into a singularity | 25 | 27 |
| It becomes a fuzzball | 15 | 16 |
| It bounces out into our universe | 11 | 12 |
| It bounces out into another universe | 12 | 13 |
| Other | 22 | 24 |
| No Opinion | 6 | 7 |
| Total | 91 | 100 |
VI.0.4 Testing Quantum Gravity Models with Black Holes
| Answer Option | Number of Votes | Percentage |
| Black hole shadows are not dark | 8 | 8 |
| Effective Field Theory corrections to inspiral/ringdown | 30 | 31 |
| Gravitational Wave Echoes | 13 | 14 |
| Deviations from Kerr Spacetime Multipoles | 13 | 14 |
| Other | 24 | 25 |
| No opinion | 8 | 8 |
| Total | 96 | 100 |
VI.0.5 Interpretation of Quantum Mechanics
| Answer Option | Number of Votes | Percentage |
| Copenhagen | 24 | 28 |
| Many Worlds | 11 | 13 |
| Qbism | 0 | 0 |
| Pilot Wave/de Broglie–Bohm | 3 | 3 |
| Collapse Theories | 6 | 7 |
| Consistent Histories | 4 | 5 |
| Other | 12 | 14 |
| No Opinion | 27 | 31 |
| Total | 87 |
VI.0.6 Dark Matter
| Answer Option | Number of Votes | Percentage |
| Light particle dark matter (e.g., axions) | 12 | 13 |
| Heavy particle dark matter (e.g., WIMPs) | 9 | 10 |
| Some form of modified gravity | 3 | 3 |
| Primordial black holes | 14 | 16 |
| Some hybrid | 23 | 26 |
| Effect of quantum gravity | 6 | 7 |
| Other | 12 | 13 |
| No opinion | 11 | 12 |
| Total | 90 | 100 |
VI.0.7 Early Universe Cosmology
| Answer Option | Number of Votes | Percentage |
| Cosmic Inflation | 37 | 40 |
| Bouncing or Cyclic Universe | 9 | 10 |
| Some quantum gravity model | 15 | 16 |
| Breakdown of Lorentz Invariance | 5 | 5 |
| Not inflation but don’t have an alternative | 10 | 11 |
| Other | 8 | 9 |
| No Opinion | 9 | 10 |
| Total | 93 | 100 |
VI.0.8 Meaning of the Big Bang
| Answer Option | Number of Votes | Percentage |
| An absolute beginning of time with a singularity at its start | 10 | 12 |
| An absolute beginning of time but without a singularity | 5 | 6 |
| A theory that says the universe evolved from a hot dense state that says nothing about whether there was an absolute beginning of time or not | 55 | 67 |
| Other | 8 | |
| No opinion | 4 | |
| Total | 82 |
VI.0.9 Anthropic Coincidences and Fine tuning
| Answer Option | Number of Votes | Percentage |
| The values of the constants are set by necessity | 8 | 10 |
| They are brute facts or need no further explanation | 26 | 32 |
| Explained by anthropic selection in a multiverse | 20 | 25 |
| Explained by a Darwinian process occurring in the cosmos | 8 | 10 |
| Explained by an intelligent designer | 2 | 2 |
| Other | 7 | 9 |
| No opinion | 10 | 12 |
| Total | 81 | 100 |
VI.0.10 Dark Energy
| Answer Option | Number of Votes | Percentage |
| A cosmological constant | 32 | 37 |
| A time varying field | 13 | 15 |
| An effect of quantum gravity | 12 | 14 |
| Some other modification to gravity | 15 | 17 |
| Other | 6 | 7 |
| No option | 8 | 9 |
| Total | 86 | 100 |
VI.0.11 Hubble Tension
| Answer Option | Number of Votes | Percentage |
| Systematic Error in Supernova Data | 27 | 35 |
| Systematic Error in CMB/Baryonic Acoustic Oscillation Data | 4 | 5 |
| Early Dark Energy | 5 | 6 |
| Other New Particles or Fields | 1 | 1 |
| Modified or Quantum Fields | 12 | 16 |
| Other | 10 | 13 |
| No opinion | 18 | 23 |
| Total | 77 |